OpenStudy (anonymous):

how do i solve : G= (x2-8x+16) how do i solve : G= (x2-8x+16) @Mathematics

OpenStudy (anonymous):

is that supposed to be x squared (x-4)^2

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[G=x{^2}-8x+16\] Correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2 - 8x + 16 x^2 - 4x - 4x + 16 x(x - 4) -4(x - 4) (x - 4)(x - 4) so G = (x - 4)(x - 4)

OpenStudy (anonymous):

Thank you ! :) but then how did i get (x-7) and (x-1) ?

OpenStudy (anonymous):

Maybe you factorised it wrong....☺ If u send the full solution that you made, I can tell u where u went wrong.....

OpenStudy (anonymous):

:) thankss agaain ! G= (x^-8x+16) = (x-4)-9 = (x-4-3) (x-4+3) = (x-7) (x-1)

OpenStudy (anonymous):

How did u get (x -4)-9 in the second step ??? .....this is wrong !!! what is yr logic for that?? pls explain in words..

OpenStudy (anonymous):

:P !! aha well this is how my "teacher" solved it !

OpenStudy (anonymous):

it is wrong !!! Is the "teacher" your school teacher or some personal tutor or some online teacher ???

OpenStudy (anonymous):

NO ! schooool teacher imagine that !

OpenStudy (anonymous):

are u frm USA ???

OpenStudy (anonymous):

:P no im lebanese

OpenStudy (anonymous):

and yourself ? :)

OpenStudy (anonymous):

I am an Indian..☺

OpenStudy (anonymous):

coool !my mom is indian too ! :)

OpenStudy (anonymous):

Really !!!! What's her name n what part of India is she from???

OpenStudy (anonymous):

She`s canadian indian :P

OpenStudy (anonymous):

and do you mind telling me where i went wrong in the math question :P

OpenStudy (anonymous):

i have a test coming up soon and im struggling alot with this subject

OpenStudy (anonymous):

do u hv a SKYPE account??

OpenStudy (anonymous):

mmm nope why :P ?

OpenStudy (anonymous):

Because on SKYPE we can have live voice chat and also share a whiteboard and explaining will be faster and better

OpenStudy (anonymous):

that`st true , but i would just like to discuss it with my parents if they are ok with it

OpenStudy (anonymous):

will that take time??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok I will try to explain here......

OpenStudy (anonymous):

cool , try your best if you don`t mind and i appreciate your help alot , thanks !

OpenStudy (anonymous):

This is a quadratic equation.......its general form is ax^2 + bx + c where a=coefficient of x^2 term , b=coefficient of x term and c=constant So first we put the given equation in general form x^2 + (- 8)x + 16, so here a = 1, b= -8 and c = 16 Now we find the product a*c = 1*16 = 16 Then we have to find two factors of 16, say P and Q, so that P*Q=a*c and P+Q=b Is it clear till now???

OpenStudy (anonymous):

yes it is :) thank you very much !

OpenStudy (anonymous):

Ok, now we can factorise 16 as 16 = 1 * 16 or -1 * -16 = 2 * 8 or -2 * -8 = 4 * 4 or -4 * -4 = 8 * 2 or -8 * -2 = 16 * 1 or -16 * -1 (notice that the last two factors are same combination as first two !!! Out of these combinations we find that only one combination fulfills both the requirements "Then we have to find two factors of 16, say P and Q, so that P*Q=a*c and P+Q=b" -4 * -4 = 16 (a*c) and (-4) + (-4) = -8 (b) so we split the middle term bx as follows : -8x = -4x - 4x x^2 - 8x + 16 x^2 - 4x - 4x + 16 Now we take out the common factor of the first two terms and then of the last two terms as follows to get: x(x - 4) -4(x - 4) (x - 4)(x - 4) this way we have factorised.........

OpenStudy (anonymous):

WOOOOOOOOOOOOOOOW ! that`s amazing ! thank you soooooooo muchh ! you helped me alot my friend :) i do appreciate it alot , 2morow ill be online the same time or earlier , maybe we can talk tomorow , i have to sleeep now , so takecare and goodnight !

OpenStudy (anonymous):

best of luck and Good Night !!!

OpenStudy (anonymous):

thanks youu too night !